# The Converse Madelung Question Schrödinger Equation from Minimal Axioms

### v1

@@ Appendix G Appendix: Projective Superposition Stress-Test
constitutes an independent falsifier of non-Fisher dynamics. This closes the experimental triangle between reversibility, linear superposition, and Fisher curvature.
@@ Appendix I Appendix: Bargmann Central Extension and Mass Superselection
Operational falsifier.
@@
We define $P_{i}=\int\rho\,\partial_{i}S\,dx$ and $K_{i}=m\int\rho\,x_{i}\,dx-tP_{i}$. With the canonical bracket $\{S(x),\rho(y)\}=-\delta(x-y)$ (so $\{F,G\}=\int(\delta F/\delta\rho\,\delta G/\delta S-\delta F/\delta S\,\delta G/\delta\rho)\,dx$) and fields in the boundary classes of Appendix A (so surface terms vanish), the generators satisfy
@@
Here $\{H,P_{i}\}=0$ holds provided $V$ has no explicit spatial dependence, all integrations by parts use the boundary classes of Appendix A, and $K_{i}$ is well defined under the finite first moment condition $\int(1+|x|)\rho\,dx<\infty$.
@@
The $t$-dependent terms are $t\int(\partial_{t}\rho\,\partial_{j}S-\partial_{t}S\,\partial_{j}\rho)dx=t\,\frac{d}{dt}\!\int\rho\,\partial_{j}S\,dx=t\,\frac{dP_{j}}{dt}$. This vanishes for a translation-invariant potential, as $\{P_{j},H\}=0$. The remaining term is:
@@
Integration by parts (vanishing boundary flux) yields:
@@
The antisymmetric $t$-term is a total divergence and integrates to zero under the boundary classes. Integrating the last term by parts gives
@@
Thus $\{H,K\}=-P$, while $\{H,P\}=0$ from translational invariance. The remaining brackets $\{P_{i},P_{j}\}=0$ and $\{K_{i},P_{j}\}=-\,\{P_{j},K_{i}\}=m\,\delta_{ij}\!\int\!\rho\,dx$ follow immediately, realising the Bargmann (Galilean) algebra with central charge $m$; under $\int\rho\,dx=1$ this is $\{K_{i},P_{j}\}=m\,\delta_{ij}$. Hence the canonical bracket and Fisher-regularised Hamiltonian are exactly Galilean covariant under the stated conditions.
@@
with the Fisher coefficient fixed by $\alpha=\kappa^{2}/(2m)$.
@@
Linearity of the target PDE forbids any quadratic or higher dependence on the state variables beyond what occurs inside $\psi$ and $\nabla\psi$. The only vector available at first order is $\nabla S$. Thus $\nabla G$ must be proportional to $\nabla S$ with a state-independent proportionality. Since $G$ is local and gauge-covariant, this implies
@@
Locality and separability of coordinate directions are essential. Any attempt to repair a mismatch by particle-dependent rephasings would break gauge covariance and the single complex structure on configuration space. Exchange symmetry is imposed at the level of the state space and does not affect the argument.
@@
For completeness we reproduce the entropy calculation in full detail. Let the continuity law include a diffusion term $\partial_{t}\rho=-\nabla\!\cdot(\rho v)+D\Delta\rho$, $v=\nabla S/m$, $D\in\mathbb{R}$. The Shannon entropy $S_{\mathrm{Sh}}[\rho]=\int\rho\ln\rho\,dx$ satisfies
$\frac{dS_{\mathrm{Sh}}}{dt}$ $=\int(1+\ln\rho)\,\rho_{t}\,dx=-\int(1+\ln\rho)\,\nabla\!\cdot(\rho v)\,dx+D\int(1+\ln\rho)\,\Delta\rho\,dx$
$=\int\rho\,v\cdot\nabla(\ln\rho)\,dx-D\int\nabla(\ln\rho)\cdot\nabla\rho\,dx$
$=\int\rho\,\nabla\!\cdot v\,dx+D\int\frac{|\nabla\rho|^{2}}{\rho}\,dx.$
Under time reversal $t\!\mapsto\!-t$, $v\!\mapsto\!-v$, the first term flips sign while the second does not. Hence invariance requires $D=0$, isolating the reversible corner. All boundary integrals vanish under the classes of Appendix A, ensuring mathematical closure.
@@
$\mathcal{R}_{\mathrm{cont}}=\frac{\langle\,|\rho_{t}+\nabla\!\cdot(\rho v)|^{2}\,\rangle}{\langle\,|\rho_{t}|^{2}+|\nabla\!\cdot(\rho v)|^{2}\,\rangle},\qquad\mathcal{R}_{\mathrm{HJ}}(\alpha)=\frac{\left\langle\,\left|S_{t}+\frac{|\nabla S|^{2}}{2m}+V-Q_{\alpha}\right|^{2}\right\rangle}{\left\langle\,|S_{t}|^{2}+\left|\frac{|\nabla S|^{2}}{2m}+V\right|^{2}+|Q_{\alpha}|^{2}\right\rangle}$
@@
which is the positive sign used in the body text and here.
@@ Acknowledgements
We thank acquaintances for valuable discussions, feedback and cross-checks. No external funding was received. All work was produced independently without affiliation. Author’s email is contact@nomogenetics.com.

### v2

@@ Appendix G Appendix: Projective Superposition Stress-Test
This constitutes an independent falsifier of non-Fisher dynamics within our admissible class. This closes the experimental triangle between reversibility, linear superposition, and Fisher curvature.
@@ Appendix I Appendix: Bargmann Central Extension and Mass Superselection
Numerical falsifier of admissibility within this class
@@
In this appendix we work on either the torus $T^{d}$, on $\mathbb{R}^{d}$ with sufficient decay at infinity, or on bounded domains with Dirichlet data $R|_{\partial\Omega}=0$. In these cases all surface terms in the generator algebra vanish, so the Bargmann relations can be checked directly. On bounded domains with Neumann data the reflecting walls break global Galilean boost symmetry and the boost generators acquire boundary contributions, which we do not need for the classification developed in the main text.
We define $P_{i}=\int\rho\,\partial_{i}S\,dx$ and $K_{i}=m\int\rho\,x_{i}\,dx-tP_{i}$. With the canonical bracket $\{S(x),\rho(y)\}=-\delta(x-y)$ (so $\{F,G\}=\int(\delta F/\delta\rho\,\delta G/\delta S-\delta F/\delta S\,\delta G/\delta\rho)\,dx$) and fields satisfying the periodic, decaying, or Dirichlet boundary conditions of Appendix A (so surface terms vanish), the generators satisfy
@@
Here $\{H,P_{i}\}=0$ holds provided $V$ has no explicit spatial dependence, all integrations by parts use the periodic, decaying, or Dirichlet boundary conditions of Appendix A, and $K_{i}$ is well defined under the finite first moment condition $\int(1+|x|)\rho\,dx<\infty$.
@@
The $t$-dependent terms are
$t\int(\partial_{t}\rho\,\partial_{j}S-\partial_{t}S\,\partial_{j}\rho)dx=t\,\frac{d}{dt}\!\int\rho\,\partial_{j}S\,dx=t\,\frac{dP_{j}}{dt}.$
This vanishes for a translation-invariant potential, as $\{P_{j},H\}=0$. The remaining term is:
@@
Integration by parts (vanishing boundary flux under the periodic, decaying, or Dirichlet boundary conditions) yields:
@@
The antisymmetric $t$-term is a total divergence and integrates to zero under the periodic, decaying, or Dirichlet boundary conditions. Integrating the last term by parts gives
@@
Thus $\{H,K\}=-P$, while $\{H,P\}=0$ from translational invariance. The remaining brackets $\{P_{i},P_{j}\}=0$ and $\{K_{i},P_{j}\}=-\,\{P_{j},K_{i}\}=m\,\delta_{ij}\!\int\!\rho\,dx$ follow immediately, realising the Bargmann (Galilean) algebra with central charge $m$; under $\int\rho\,dx=1$ this is $\{K_{i},P_{j}\}=m\,\delta_{ij}$. Hence the canonical bracket and Fisher-regularised Hamiltonian are exactly Galilean covariant in these settings.
@@
with the Fisher coefficient fixed by $\alpha=\kappa^{2}/(2m)$. Identifying $\kappa=\hbar$ then yields the Fisher scale $\alpha=\hbar^{2}/(2m)$ used in the main text.
@@
Linearity and locality of the target PDE forbid any explicit quadratic or higher dependence on $(\rho,S)$ beyond what is already encoded inside $\psi$ and its first derivatives.
The only vector available at first order is $\nabla S$. Thus $\nabla G$ must be proportional to $\nabla S$ with a state-independent proportionality. Since $G$ is local and gauge-covariant, this implies
@@
Taken together with the curvature classification in the main text, this shows that there is no second local change of variables hiding in the background that could secretly linearise a different reversible theory while still respecting all of the axioms.
@@
Locality and separability of coordinate directions are essential. Any attempt to repair a mismatch by particle-dependent rephasings would necessarily destroy gauge covariance and the assumption of a single local complex structure on configuration space. Exchange symmetry is imposed at the level of the state space and does not affect the argument.
@@
For completeness we reproduce the entropy calculation in full detail. Let the continuity law include a diffusion term
$\partial_{t}\rho\;=\;-\,\nabla\!\cdot(\rho v)\;+\;D\,\Delta\rho,\qquad v=\nabla S/m,\quad D\in\mathbb{R}.$
We take the Shannon entropy to be
$S_{\mathrm{Sh}}[\rho]\;:=\;-\int_{\Omega}\rho\ln\rho\,dx,$
which is minus the Boltzmann $H$-functional. Then
$\frac{dS_{\mathrm{Sh}}}{dt}$ $=-\int_{\Omega}(1+\ln\rho)\,\partial_{t}\rho\,dx$
$=-\int_{\Omega}(1+\ln\rho)\,\big(-\nabla\!\cdot(\rho v)+D\Delta\rho\big)\,dx$
$=\int_{\Omega}(1+\ln\rho)\,\nabla\!\cdot(\rho v)\,dx\;-\;D\int_{\Omega}(1+\ln\rho)\,\Delta\rho\,dx$
$=-\int_{\Omega}\rho\,v\cdot\nabla(\ln\rho)\,dx\;+\;D\int_{\Omega}\nabla(\ln\rho)\cdot\nabla\rho\,dx$
$=-\int_{\Omega}v\cdot\nabla\rho\,dx\;+\;D\int_{\Omega}\frac{|\nabla\rho|^{2}}{\rho}\,dx$
$=\int_{\Omega}\rho\,\nabla\!\cdot v\,dx\;+\;D\int_{\Omega}\frac{|\nabla\rho|^{2}}{\rho}\,dx.$
Under time reversal $t\mapsto-t$, $v\mapsto-v$, the first term flips sign while the second does not. Hence exact time-reversal invariance of $S_{\mathrm{Sh}}$ requires $D=0$, isolating the reversible corner. All boundary integrals vanish under the classes of Appendix A, ensuring mathematical closure. In particular, for $D\neq 0$ the sign-definite Fisher term cannot be cancelled by any choice of velocity field $v$, so $dS_{\mathrm{Sh}}/dt$ fails to reverse under $t\mapsto-t$.
@@
$\mathcal{R}_{\mathrm{cont}}=\frac{\langle\,|\rho_{t}+\nabla\!\cdot(\rho v)|^{2}\,\rangle}{\langle\,|\rho_{t}|^{2}+|\nabla\!\cdot(\rho v)|^{2}\,\rangle},\qquad\mathcal{R}_{\mathrm{HJ}}(\alpha)=\frac{\left\langle\,\left|S_{t}+\frac{|\nabla S|^{2}}{2m}+V+Q_{\alpha}\right|^{2}\right\rangle}{\left\langle\,|S_{t}|^{2}+\left|\frac{|\nabla S|^{2}}{2m}+V\right|^{2}+|Q_{\alpha}|^{2}\right\rangle}$
@@
which fixes the positive sign convention for $\partial_{t}\rho$ used throughout the body text and here.
@@ Acknowledgements
We thank colleagues and acquaintances for valuable discussions, feedback and cross-checks. No external funding was received. All work was produced independently without affiliation. Author’s email is contact@nomogenetics.com.